OpenStudy (anonymous):

What is the value of the missing exponent? (Type only the numerical value of the missing exponent. Do not skip any spaces.) (4x4)?=16x8

OpenStudy (paxpolaris):

\[\Large {\left( 4x^4 \right)^2\\ = 4^2\left( x^4 \right)^2\\ =16\cdot x ^{(4\times2)}\\=16x^8}\]

OpenStudy (anonymous):

so the missing exponent is 8?

OpenStudy (anonymous):

huh?@ paxpolaris

OpenStudy (anonymous):

@PaxPolaris

OpenStudy (anonymous):

hello??

ganeshie8 (ganeshie8):

no

ganeshie8 (ganeshie8):

continuing after paxpolaris steps, you should get exponent = 2

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

np i hope u get hw thats the case :)